OpenStudy (anonymous):

what is the y intercept of the line with the equation x - 3y = -12

OpenStudy (anonymous):

substitute x=0 and solve for y

jimthompson5910 (jim_thompson5910):

x - 3y = -12 0 - 3y = -12 .. plug in x = 0 -3y = -12 y = ???

OpenStudy (anonymous):

y = -4?

jimthompson5910 (jim_thompson5910):

very very close, but no

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

positive 4?

jimthompson5910 (jim_thompson5910):

much better, y = 4

jimthompson5910 (jim_thompson5910):

the y intercept is y = 4 which means that it is the point (0,4)